# Mechanical detection of nuclear decays

### v1

@@ Acknowledgments.
We thank Daniel Carney and Giorgio Gratta for useful discussions, and Fernando Monteiro and Gadi Afek for early work on the techniques used here. This work was supported through Quantum Horizons: QIS Research and Innovation for Nuclear Science Award DE-SC0023672, and in part by ONR Grant N00014-23-1-2600 and NSF Grant PHY-2109329.
@@
Commercially produced silica microspheres are used in this work (Corpuscular, C-SIO-3.0). The mean radius of the spheres measured by an optical microscope is $1.4$ $\rm{\mu}$m, with a variation of 5% between different spheres. The spheres are loaded with ${}^{212}$b using a flow-through ${}^{220}$n emanation source (Pylon TH-1025, activity of 10 kBq on Jan. 1, 2023). Dry microspheres are applied to the surface of a glass cover slip, which is positioned above a negatively biased electrode in a dedicated vacuum chamber for the implantation process. The chamber is pumped to a pressure of 10 mbar and maintained at this pressure while ${}^{220}$n is transported to the chamber by a flow of 4.6 sccm of nitrogen carrier gas. Simulations using the Stopping and Range of Ions in Matter (SRIM) software package [50, 51] indicate this pressure is sufficient to stop the majority of ${}^{216}$o recoils following decay of the ${}^{220}$n nucleus before they hit the chamber walls. Previous measurements indicate a substantial fraction of ${}^{216}$o daughters are expected to remain as singly ionized ${}^{216}$${}^{+}$ after stopping in dilute gas [52, 53]. These ions are drifted to the surface of the cover slip on which the spheres are deposited by the electric field from a 250 $\mu$m diameter pin positioned just below the opposite side of the cover slip and biased to a potential of $-1$ kV (see Fig. 6).
[…]
@@
After loading, the cover slip is positioned in the main trapping chamber above the optical trap and spheres are detached from its surface by using a piezoelectric actuator to vibrate the cover slip. Once a sphere released from the cover slip is trapped, the chamber is pumped to high vacuum and charge and recoil measurements begin. Due to the relatively short $T_{1/2}$ of ${}^{212}$b, no radioactive decay products will remain in the vacuum chamber after a few days following the loading of the cover slip.
@@
Figure 7: (left) Simulation of the implantation of ${}^{212}$b daughters into the microspheres following $\alpha$ decays of ${}^{216}$o on the sphere surface. The red lines indicate the projection into the $x$-$y$ plane of simulated trajectories of randomly oriented ${}^{212}$b daughter recoils starting from $(x,y,z)=(1.5,0,0)\ \mu$m. Approximately 50% of decays are stopped within the sphere (gray). (right) Distribution of implantation depths as a function of the nearest distance from the sphere surface (black). The median simulated implantation depth is 29 nm, with 95% of ${}^{212}$b implanted within 60 nm of the surface. The orange distribution indicates the total distance between the ${}^{216}$o decay location and final ${}^{212}$b location.
@@
The electric field, as a function of applied voltage, is calculated using a finite element analysis (FEA) simulation of the electrode configuration with the COMSOL Multiphysics software program [54]. The exact geometry of the electrodes is modeled by the simulation, including features such as the holes in the electrode faces needed to allow the trapping and readout lasers, as well as the filament and UV light source for charge control. The electric field simulated for a $+$1 V potential applied to the $+x$ electrode with all other electrodes at ground is shown in Fig. 8. Using these simulations, the total impulse applied is determined after integrating the voltage waveform measured at the monitor output of the high voltage supply (Tabor 9400A). For each calibration, the sphere’s motion in the $x$, $y$, and $z$ direction is measured while a waveform consisting of repeated square pulses with a pulse width of 1 ms and repetition rate of 2.3 Hz is applied. The voltage is set to provide a total impulse of approximately 200 MeV/c, i.e., typically a 6.7 V/cm electric field is applied during the pulse, with a net charge on the sphere set to be $|Q|=1\ e$. A total of 300 s of data are recorded during each calibration, yielding approximately 700 pulses per calibration. The calibration procedure is repeated for pulses applied in the $x$, $y$, and $z$ directions individually.
@@
[…]
@@
Figure 10: Example of position ASD for the $x$ (left), $y$ (center), and $z$ (right) directions measured by averaging 500 s of noise data. The colored lines correspond to the expected noise for a simple harmonic oscillator with the resonator parameters shown in Table 1 driven by white noise.
@@
In addition to the stationary noise whose ASD is shown in Fig. 10, random pulses occur throughout the data taking period that are consistent with transient impulses or other force noise that is not correlated with a decay within the sphere. In past work, such impulses were seen to arise from vibrationally or acoustically transmitted noise sources that exert a force on the sensor [27]. An example is shown in Fig. 11. Here, in addition to the impulse seen around 75 s (a candidate ${}^{212}$i $\alpha$ decay), which is correlated with a sphere charge change, a large signal consistent with the expected pulse shape for an impulse in the $x$ direction is seen around 50 s, but with no charge change. In addition, noise bursts inconsistent with the impulse response are seen in the $y$ and $z$ channels around 18 s and 65 s. No impulse is seen at the time of the observed charge change around 97 s (consistent with the undetectable momentum of a candidate $\beta$ decay from ${}^{208}$l).
@@
Figure 11: Example event containing candidate ${}^{212}$i–${}^{208}$l coincident decays, in addition to transient noise that is not correlated with a charge change of the sphere. The upper three panels show the reconstructed impulse amplitude versus time in the $x$, $y$, and $z$, directions, while the lower panel shows the reconstructed charge. Charge changes observed around 75 s and 97 s are highlighted by the blue shaded regions.
@@
As shown in Fig. 7, we expect the activity loaded within the sphere will be implanted within $\lesssim 60$ nm of the sphere surface. Lower chain $\alpha$ decays following the initial ${}^{212}$b implantation can thus in some cases eject the daughter nucleus from the surface of the sphere (which may additionally eject secondary O or Si nuclei). In addition, because of the poorer momentum resolution in the $y$ and $z$ directions (and uncertainty due to time-dependent drifts in the calibration for these directions), in this work, we focus on accurately reconstructing only the portion of the momentum for each decay the $x$ direction. A Monte Carlo simulation is used to both account for the escape of nuclear recoils throughout the decay chain and determine the expected projected momentum in $x$ for each simulated decay.
@@
The momentum spectrum used to fit the observed data in Fig. 4, was calculated using a simulation of the full decay chain based on SRIM. This simulation was used to produce $10^{7}$ decays starting from ${}^{216}$o deposited at a random location on the surface of the sphere. This simulation makes use of the half-life, recoil energy, and branching ratio data from the NuDat 3.0 database [32]. For each $\alpha$ decay in the chain, a trajectory is first simulated with SRIM, and custom code is used to follow this trajectory through the geometry of the sphere until the nuclear recoil stops within the sphere or exits its surface. If the nuclear recoil exits the sphere, its remaining momentum upon exiting the sphere is recorded and the simulation terminates for that event. If the nuclear recoil stops, subsequent decays are then modeled until they either reach stable ${}^{208}$b at the end of the decay chain or the subsequent daughters exit the sphere. Examples of simulated events using this code are shown in Fig. 14. SRIM simulations also indicate the scattering of the emitted $\alpha$ particles within the micron-sized spheres is negligible, and for simplicity the $\alpha$ is assumed to exit the sphere with momentum equal to its momentum at the decay location, with only the nuclear recoils of the $\alpha$ decay daughters simulated in detail.
@@
Figure 14: Example simulated decay events for the full decay chain starting from a randomly oriented $\alpha$ decay of ${}^{216}$o at the sphere surface. The upper row of plots shows an example event with a ${}^{212}$i $\alpha$ decay, which reaches stable ${}^{208}$b without the nuclear recoil exiting the sphere. The lower plots show a decay chain with a ${}^{212}$o $\alpha$ decay, for which the daughter ${}^{208}$b nuclear recoil exits the sphere. For each event, the left plot shows the 3D trajectory, with the projected paths in the $x$-$y$ and $y$-$z$ planes shown in the middle panels. The gray line indicates the edge of the sphere (with radius $1.5\ \mu$m). The right panel shows the radius of the daughter as a function of its energy loss throughout the trajectory. Daughters of $\alpha$ decays are shown as large markers. Nuclear recoils from $\beta$ decays (small markers) will not travel significant distance from the decay location and their trajectories are not simulated in detail.
@@
The total momentum transferred to the sphere in each decay is calculated in the simulation using momentum conservation as $\vec{p}_{sph}=-(\vec{p}_{\alpha}+\vec{p}_{NR})$, where $\vec{p}_{\alpha}$ is the momentum of the $\alpha$ particle and $\vec{p}_{NR}$ is the momentum of the nuclear recoil simulated by SRIM for each decay upon exiting the sphere (if the nuclear recoil stops within the sphere, $\vec{p}_{NR}=0$). Each component of $\vec{p}_{sph}$ is then smeared by a noise value randomly selected from the resolution measured in the in situ calibrations for each sphere, weighted by live time. The resulting momentum distribution is shown in Fig. 15 (left), for an example where we assume the noise in all three degrees of freedom is improved to the currently measured noise in the $x$ direction. The low momentum tail of the distribution corresponds to events in which the daughter nuclear recoil escapes before fully transferring its momentum to the sphere, due to the shallow implantation depth of the ${}^{212}$b in this work. Future work will aim to characterize the escape of these daughters, and possibly veto them based on the corresponding charge change. Alternative sphere synthesis methods could also allow the portion of the sphere containing the activity to be capped by a $\gtrsim$100 nm thick inactive silica layer, preventing escape of these daughter recoils [31].
@@
Figure 15: (left) Simulation of the magnitude of the momentum transferred to the sphere from the Monte Carlo described in the text, assuming an equal momentum resolution independent of the decay direction. The peak in the momentum distribution corresponds to ${}^{212}$i (dark orange) and ${}^{212}$o (light orange) $\alpha$ decays for which the nuclear recoil stops in the sphere and its full momentum is transferred to the sphere’s center-of-mass motion. Due to the shallow implantation of the ${}^{212}$b parent, approximately 25% of decays have reduced momentum due to escape of the nuclear recoil from the sphere (providing the low momentum tail below the peak). (right) Projection of the $x$ component of the simulated momentum for each event.

### v2

@@ Acknowledgments.
We thank Daniel Carney and Giorgio Gratta for useful discussions, and Fernando Monteiro and Gadi Afek for early work on the techniques used here. This work was supported through the DOE Office of Nuclear Physics, Quantum Horizons Award DE-SC0023672, and in part by ONR Grant N00014-23-1-2600 and NSF Grant PHY-2109329.
@@
Commercially produced silica microspheres are used in this work (Corpuscular, C-SIO-3.0). The mean radius of the spheres measured by an optical microscope is $1.4$ $\rm{\mu}$m, with a variation of 5% between different spheres. The spheres are loaded with 212Pb using a flow-through 220Rn emanation source (Pylon TH-1025, activity of 10 kBq on Jan. 1, 2023). Dry microspheres are applied to the surface of a glass cover slip, which is positioned above a negatively biased electrode in a dedicated vacuum chamber for the implantation process. The chamber is pumped to a pressure of 10 mbar and maintained at this pressure while 220Rn is transported to the chamber by a flow of 4.6 sccm of nitrogen carrier gas. Simulations using the Stopping and Range of Ions in Matter (SRIM) software package [51, 52] indicate this pressure is sufficient to stop the majority of 216Po recoils following decay of the 220Rn nucleus before they hit the chamber walls. Previous measurements indicate a substantial fraction of 216Po daughters are expected to remain as singly ionized 216Po+ after stopping in dilute gas [53, 54]. These ions are drifted to the surface of the cover slip on which the spheres are deposited by the electric field from a 250 $\mu$m diameter pin positioned just below the opposite side of the cover slip and biased to a potential of $-1$ kV (see Fig. 6).
Figure 6: (a, left) Schematic of the implantation chamber. A metal pin is placed at the bottom center of the chamber to apply the DC high voltage. A Delrin disk surrounds the pin, securing the sphere-loaded glass cover slip and stabilizing the pin’s position. Two ports on the chamber’s top lid allow gas inflow and outflow. The inset provides a magnified view of the spheres on the upper surface of the cover slip, with the pin positioned underneath. (b, right) Cross-section of a simulation depicting the electric field lines within the implantation chamber, illustrating the percentage of the volume directed to the approximately 1 mm2 region containing the spheres when the pin is biased to a voltage of $-1$ kV. The gray region is an aluminum spacer, which, along with the chamber walls, is grounded. The yellow region is the Delrin disk, which provides mechanical support and electrical insulation of the pin.
The spheres are loaded with 220Rn daughters for a total period several times longer than the 212Pb half-life, so that secular equilibrium is reached and the maximum sphere activity is achieved (typically 1-2 days). The typical measured sphere activity following loading translates to a average activity of 200 Bq/mm2 of 216Po+ reaching the surface of the cover slip (translating to approximately 1 mBq activity per sphere). The 216Po+ ions drifted to the sphere surface promptly decay and, as shown in Fig. 7, approximately 50% of the time these decays implant the 212Pb daughter into the surface of the sphere at a depth up to 60 nm.
@@
After loading, the cover slip is positioned in the main trapping chamber above the optical trap and spheres are detached from its surface by using a piezoelectric actuator to vibrate the cover slip. Once a sphere released from the cover slip is trapped, the chamber is pumped to high vacuum and charge and recoil measurements begin. Due to the relatively short $T_{1/2}$ of 212Pb, no radioactive decay products will remain in the vacuum chamber after a few days following the loading of the cover slip.
@@
Figure 7: (left) Simulation of the implantation of 212Pb daughters into the microspheres following $\alpha$ decays of 216Po on the sphere surface. The red lines indicate the projection into the $x$-$y$ plane of simulated trajectories of randomly oriented 212Pb daughter recoils starting from $(x,y,z)=(1.5,0,0)\ \mu$m. Approximately 50% of decays are stopped within the sphere (gray). (right) Distribution of implantation depths as a function of the nearest distance from the sphere surface (black). The median simulated implantation depth is 29 nm, with 95% of 212Pb implanted within 60 nm of the surface. The orange distribution indicates the total distance between the 216Po decay location and final 212Pb location.
@@
The electric field, as a function of applied voltage, is calculated using a finite element analysis (FEA) simulation of the electrode configuration with the COMSOL Multiphysics software program [55]. The exact geometry of the electrodes is modeled by the simulation, including features such as the holes in the electrode faces needed to allow the trapping and readout lasers, as well as the filament and UV light source for charge control. The electric field simulated for a $+$1 V potential applied to the $+x$ electrode with all other electrodes at ground is shown in Fig. 8. Using these simulations, the total impulse applied is determined after integrating the voltage waveform measured at the monitor output of the high voltage supply (Tabor 9400A). For each calibration, the sphere’s motion in the $x$, $y$, and $z$ direction is measured while a waveform consisting of repeated square pulses with a pulse width of 1 ms and repetition rate of 2.3 Hz is applied. The voltage is set to provide a total impulse of approximately 200 MeV/c(comparable to the impulse during an $\alpha$ decay), i.e., typically a 6.7 V/cm electric field is applied during the pulse, with a net charge on the sphere set to be $|Q|=1\ e$. A total of 300 s of data are recorded during each calibration, yielding approximately 700 pulses per calibration. The calibration procedure is repeated for pulses applied in the $x$, $y$, and $z$ directions individually.
@@
Finally, there is uncertainty in the sphere mass since the radius and density are not precisely measured for each sphere, and previous studies have found lower sphere densities than expected for pure fused silica [56]. While an advantage of the calibration method described above is that this uncertainty does not affect the impulse calibration, it may affect the calibration of position that appears for reference in figures throughout this work. For the position calibration, we assume a nominal density of 2.0 g/cm3 [57] for the sphere and a radius of 1.4 $\mu$m, leading to an estimated sphere mass of 23 pg. We estimate that there is a $\pm$30systematic uncertainty on the calibrated position response for a given sphere due to the uncertainty on its mass.
@@
Figure 10: Example of position ASD for the $x$ (left), $y$ (center), and $z$ (right) directions measured by averaging 500 s of noise data. The colored lines correspond to the expected noise for a simple harmonic oscillator with the resonator parameters shown in Table 1 driven by white noise. The displacement sensitivity of this system is $\lesssim 10^{-11}$ m/$\sqrt{\mathrm{Hz}}$.
@@
In addition to the stationary noise whose ASD is shown in Fig. 10, random pulses occur throughout the data taking period that are consistent with transient impulses or other force noise that is not correlated with a decay within the sphere. In past work, such impulses were seen to arise from vibrationally or acoustically transmitted noise sources that exert a force on the sensor [27]. An example is shown in Fig. 11. Here, in addition to the impulse seen around 75 s (a candidate 212Bi $\alpha$ decay), which is correlated with a sphere charge change, a large signal consistent with the expected pulse shape for an impulse in the $x$ direction is seen around 50 s, but with no charge change. In addition, noise bursts inconsistent with the impulse response are seen in the $y$ and $z$ channels around 18 s and 65 s. No impulse is seen at the time of the observed charge change around 97 s (consistent with the undetectable momentum of a candidate $\beta$ decay from 208Tl).
@@
Figure 11: Example event containing candidate 212Bi–208Tl coincident decays, in addition to transient noise that is not correlated with a charge change of the sphere. The upper three panels show the reconstructed impulse amplitude versus time in the $x$, $y$, and $z$, directions, while the lower panel shows the reconstructed charge. Charge changes observed around 75 s and 97 s are highlighted by the blue shaded regions.
@@
As shown in Fig. 7, we expect the activity loaded within the sphere will be implanted within $\lesssim 60$ nm of the sphere surface. Lower chain $\alpha$ decays following the initial 212Pb implantation can thus in some cases eject the daughter nucleus from the surface of the sphere (which may additionally eject secondary O or Si nuclei). In addition, because of the poorer momentum resolution in the $y$ and $z$ directions (and uncertainty due to time-dependent drifts in the calibration for these directions), in this work, we focus on accurately reconstructing only the portion of the momentum for each decay the $x$ direction. A Monte Carlo simulation is used to both account for the escape of nuclear recoils throughout the decay chain and determine the expected projected momentum in $x$ for each simulated decay.
@@
The momentum spectrum used to fit the observed data in Fig. 4, was calculated using a simulation of the full decay chain based on SRIM. This simulation was used to produce $10^{7}$ decays starting from 216Po deposited at a random location on the surface of the sphere. This simulation makes use of the half-life, recoil energy, and branching ratio data from the NuDat 3.0 database [32]. For each $\alpha$ decay in the chain, a trajectory is first simulated with SRIM, and custom code is used to follow this trajectory through the geometry of the sphere until the nuclear recoil stops within the sphere or exits its surface. If the nuclear recoil exits the sphere, its remaining momentum upon exiting the sphere is recorded and the simulation terminates for that event. If the nuclear recoil stops, subsequent decays are then modeled until they either reach stable 208Pb at the end of the decay chain or the subsequent daughters exit the sphere. Examples of simulated events using this code are shown in Fig. 14. The loss of energy and momentum by the $\alpha$ within the sphere was also simulated in SRIM. For the worst case of transiting through the entire 3 $\mu$m thickness of the sphere, the energy loss within the sphere is $<350$ keV from the SRIM simulation, which corresponds to a maximum change in the momentum of the exiting of $<$5 MeV/c. Since even this maximum change is small compared to the resolution, and the median decay (after averaging over all decay angles) will have a momentum loss $<$1 MeV/c, scattering within the sphere is neglected in the simulation. For simplicity the $\alpha$ is assumed to exit the sphere with momentum equal to its momentum at the decay location, with only the nuclear recoils of the decay daughters simulated in detail.
@@
Figure 14: Example simulated decay events for the full decay chain starting from a randomly oriented $\alpha$ decay of 216Po at the sphere surface. The upper row of plots shows an example event with a 212Bi $\alpha$ decay, which reaches stable 208Pb without the nuclear recoil exiting the sphere. The lower plots show a decay chain with a 212Po $\alpha$ decay, for which the daughter 208Pb nuclear recoil exits the sphere. For each event, the left plot shows the 3D trajectory, with the projected paths in the $x$-$y$ and $y$-$z$ planes shown in the middle panels. The gray line indicates the edge of the sphere (with radius $1.5\ \mu$m). The right panel shows the radius of the daughter as a function of its energy loss throughout the trajectory. Daughters of $\alpha$ decays are shown as large markers. Nuclear recoils from $\beta$ decays (small markers) will not travel significant distance from the decay location and their trajectories are not simulated in detail.
@@
The total momentum transferred to the sphere in each decay is calculated in the simulation using momentum conservation as $\vec{p}_{sph}=-(\vec{p}_{\alpha}+\vec{p}_{NR})$, where $\vec{p}_{\alpha}$ is the momentum of the $\alpha$ particle and $\vec{p}_{NR}$ is the momentum of the nuclear recoil simulated by SRIM for each decay upon exiting the sphere (if the nuclear recoil stops within the sphere, $\vec{p}_{NR}=0$). Each component of $\vec{p}_{sph}$ is then smeared by a noise value randomly selected from the resolution measured in the in situ calibrations for each sphere, weighted by live time. The resulting momentum distribution is shown in Fig. 15 (left), for an example where we assume the noise in all three degrees of freedom is improved to the currently measured noise in the $x$ direction. The low momentum tail of the distribution corresponds to events in which the daughter nuclear recoil escapes before fully transferring its momentum to the sphere, due to the shallow implantation depth of the 212Pb in this work. Future work will aim to characterize the escape of these daughters, and possibly veto them based on the corresponding charge change. Alternative sphere synthesis methods could also allow the portion of the sphere containing the activity to be capped by a $\gtrsim$100 nm thick inactive silica layer, preventing escape of these daughter recoils [31].
@@
In addition, for isolated $\beta$ decays (i.e. 212Pb or 208Tl), the change in the net charge of the microsphere will still be detected, and the optimal filter will be used to reconstruct the coincident impulse amplitude. This is expected to produce a distribution of reconstructed events with an amplitude consistent with noise. To accurately model the reconstructed amplitudes for these $\beta$ decays, we sample noise data uniformly distributed throughout the data acquisition period and apply the same optimal filter based reconstruction at randomly selected times (uncorrelated with a charge change) that we apply to detect recoils in coincidence with a charge change. This empirically driven procedure accounts for any effects from transient noise, random coincidence with noise bursts, etc. that may be present in the data (see Sec. .5).
The resulting distribution of reconstructed amplitudes for these random times is used to produce the gray shaded distribution in Fig. 4. A total momentum distribution is then created by adding this to the results of the $\alpha$ particle recoil simulation. This distribution is then fit to the data as shown in Fig. 4, with the ratios of the 212Po and 212Bi fixed based on the decay chain simulation, but allowing the rate of the modelled $\beta$ decays to vary as a free parameter in the fit.
Figure 15: (left) Simulation of the magnitude of the momentum transferred to the sphere from the Monte Carlo described in the text, assuming an equal momentum resolution independent of the decay direction. The peak in the momentum distribution corresponds to 212Bi (dark orange) and 212Po (light orange) $\alpha$ decays for which the nuclear recoil stops in the sphere and its full momentum is transferred to the sphere’s center-of-mass motion. Due to the shallow implantation of the 212Pb parent, approximately 25% of decays have reduced momentum due to escape of the nuclear recoil from the sphere (providing the low momentum tail below the peak). (right) Projection of the $x$ component of the simulated momentum for each event.